# Decision-focused predictions via pessimistic bilevel optimization: complexity and algorithms

### v4

@@ 5.2 Computational results
In this section, we compare the decision-focused predictions obtained using the aforementioned methods by evaluating the regrets they achieve. To ensure the regrets are displayed on the same scale, and following the approach in Elmachtoub and Grigas, (2022), we use the normalized regret instead of reporting the direct regret (as defined in (1.1)). The normalized regret is defined as
@@ 5.2.1 Training set performance
Tables 2 and 3 show the performance of the methods on the training set for small instances of shortest paths and maximum weight bipartite matching, respectively. In these tables, we use the regret returned by SPO (model (13)) as the baseline. The subsequent columns display the percentage decrease in regret achieved by our methods. The method with the best performance in terms of normalized regret is highlighted in color.
In the shortest path instances shown in Table 2, our alternating method, either by itself or in tandem with local search, achieves the best performance in terms of regret. We also observe that the penalized method may often improve (decrease) the regret, but in some instances, it can dramatically increase the regret (black bold numbers). This occurs whenever the starting point is not valid for the slice given by the chosen penalty factor in (14), causing the solver to reject the solution. Table 3 shows that for bipartite matching there are cases where the the penalized method was the best. However, the alternating direction method obtained the overall best results.
These results indicate that the performance of the alternating method considerably depends on its starting point. In most cases, starting from the solution provided by the local search algorithm yields a better outcome than starting with SPO. However, since the time for SPO, LS, and ALT is divided within the one-hour time limit, there are cases where starting with SPO and allowing the alternating method to run for the majority of the time results in better performance. Overall, we believe these are highly encouraging results: all methods we develop here are tailored for improving the training performance in moderate running times, to which we succeded for these challenging instances.
In what follows, we discuss two remaining aspects: test set performance and scalability.
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.035 0.0% -14.3% -17.1% -42.9% -54.3%
50 2 0.5 0.171 0.0% -$2$5.1% -43.3% -43.3%
50 8 0.0 0.133 0.0% -13.5% +53.4% -19.5% -19.5%
50 8 0.5 0.381 -23.6% -45.1% +9.2% -67.5% -64.8%
50 16 0.0 1.254 -66.4% -70.6% -6.7% -82.0% -75.4%
50 16 0.5 0.99 -59.5% -64.5% +253.5% -71.2% -77.8%
100 2 0.0 0.078 0.0% 0.0% 0.0% -32.1% -32.1%
100 2 0.5 0.195 0.0% 0.0% 0.0% -37.9% -40.5%
100 8 0.0 0.399 0.0% 0.0% 0.0% -20.8% -20.8%
100 8 0.5 0.572 -29.0% -29.0% -29.0% -47.6% -44.9%
100 16 0.0 3.243 -61.2% -61.2% -23.0% -69.1% -13.8%
100 16 0.5 1.945 -15.3% -15.3% +76.1% -28.0% -5.3%
200 2 0.0 0.089 0.0% 0.0% 0.0% -20.2% -20.2%
200 2 0.5 0.232 -0.9% -0.9% -0.9% -11.2% -13.4%
200 8 0.0 0.693 -16.7% -16.7% -16.7% -25.4% -4.0%
200 8 0.5 0.621 -1.4% -1.4% -1.4% -22.1% -16.1%
200 16 0.0 2.14 -60.5% -60.5% -60.5% -61.5% -2.9%
200 16 0.5 6.194 -79.4% -79.4% -79.4% -80.4% -0.2%
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.065 0.0% -12.3% -46.2% -56.9% -56.9%
50 2 0.5 0.125 0.0% -15.2% -35.2% -46.4% -46.4%
50 8 0.0 0.344 0.0% -12.8% -24.4% -31.7% -31.7%
50 8 0.5 0.249 0.0% -15.3% -11.2% -39.4% -39.4%
50 16 0.0 0.316 -3.5% -11.4% +5.1% -52.2% -50.0%
50 16 0.5 0.344 -18.3% -32.0% -18.3% -74.1% -38.4%
100 2 0.0 0.091 0.0% -4.4% -47.3% -29.7% -29.7%
100 2 0.5 0.182 0.0% 0.0% -42.3% -27.5% -27.5%
100 8 0.0 0.391 -10.7% -10.7% -10.7% -40.2% -20.2%
100 8 0.5 0.29 0.0% -5.9% 0.0% -27.6% -27.6%
100 16 0.0 0.415 0.0% -6.7% 0.0% -19.8% -19.8%
100 16 0.5 0.536 -12.9% -12.9% -12.9% -38.6% -19.8%
200 2 0.0 0.103 0.0% 0.0% 0.0% -15.5% -15.5%
200 2 0.5 0.207 0.0% 0.0% 0.0% -11.6% -11.6%
200 8 0.0 0.361 0.0% 0.0% 0.0% -8.9% -8.9%
200 8 0.5 0.387 0.0% 0.0% 0.0% -7.5% -7.5%
200 16 0.0 0.478 -17.6% -17.6% -17.6% -28.9% -17.6%
200 16 0.5 0.589 -25.3% -25.3% -25.3% -33.8% -28.5%
@@ 5.2.2 Test set performance
In Table 4 we presents results for shortest path instances. We observe that in 14 out of 16 cases, one of the methods we propose here improves the SPO performance. The two variants of the alternating method, which was the best in training performance, has the best performance in 8 out of 18 test instances, however, it can increase the test regret in comparison to SPO in some cases. The results for bipartite matching, which we show in Table 5, are slightly different. Here, the alternating method, in its various implementations, achieves the best performance in 9 out of 18 instances –one more than in the shortest path instances. On the other hand, in 6 instances, no method was able to improve the test performance of SPO.
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.105 0.0% -2.9% -24.8% -11.4% -4.8%
50 2 0.5 0.304 0.0% -15.1% +6.9% -6.9% -7.2%
50 8 0.0 0.45 0.0% +8.2% -34.2% -52.7% -34.4%
50 8 0.5 1.056 -3.3% +8.4% +7.4% +16.5% +5.2%
50 16 0.0 3.762 -74.9% -73.1% -76.2% -72.8% -74.9%
50 16 0.5 3.504 -9.6% -7.4% +42.9% +6.2% -13.8%
100 2 0.0 0.099 0.0% 0.0% 0.0% -1.0% -1.0%
100 2 0.5 0.227 0.0% 0.0% 0.0% +7.0% +11.0%
100 8 0.0 0.776 0.0% 0.0% 0.0% +8.9% -13.9%
100 8 0.5 0.707 -3.0% -3.0% -3.0% -24.9% +35.5%
100 16 0.0 3.882 +39.6% +39.6% -12.6% -6.1% -7.5%
100 16 0.5 13.645 +13.9% +13.9% -8.2% +14.9% +5.0%
200 2 0.0 0.099 0.0% 0.0% 0.0% 0.0% -1.0%
200 2 0.5 0.238 +11.3% +11.3% +11.3% +10.9% -5.0%
200 8 0.0 0.591 -2.7% -2.7% -2.7% +1.9% +14.7%
200 8 0.5 1.023 -23.3% -23.3% -23.3% -19.8% -2.5%
200 16 0.0 1.836 +39.9% +39.9% +39.9% +27.6% +15.0%
200 16 0.5 3.384 +2.0% +2.0% +2.0% -5.8% -12.1%
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.118 0.0% +39.0% -2.5% -3.4% +14.4%
50 2 0.5 0.26 0.0% 0.0% -5.8% -4.2% -10.8%
50 8 0.0 0.349 0.0% +30.4% -6.9% -6.3% -10.3%
50 8 0.5 0.474 0.0% +6.1% +4.0% +6.5% +5.3%
50 16 0.0 0.781 -17.2% -20.5% -13.3% +15.1% -17.4%
50 16 0.5 0.596 -21.8% -9.4% -21.8% -9.9% +11.4%
100 2 0.0 0.118 0.0% +51.7% -16.1% +5.9% +5.9%
100 2 0.5 0.252 0.0% 0.0% -1.2% 0.0% -0.8%
100 8 0.0 0.403 +6.2% +6.2% +6.2% +5.2% 0.0%
100 8 0.5 0.431 0.0% -1.6% 0.0% -2.6% -2.1%
100 16 0.0 0.734 0.0% 0.1% 0.0% -18.5% -18.4%
100 16 0.5 0.455 +16.9% +16.9% +16.9% +52.1% -1.3%
200 2 0.0 0.126 0.0% 0.0% 0.0% -4.0% -3.2%
200 2 0.5 0.229 0.0% 0.0% 0.0% +3.1% +2.6%
200 8 0.0 0.368 0.0% 0.0% 0.0% +6.5% +6.5%
200 8 0.5 0.401 0.0% 0.0% 0.0% +1.5% +1.5%
200 16 0.0 0.739 -6.4% -6.4% -6.4% -19.4% +0.1%
200 16 0.5 0.455 +51.2% +51.2% +51.2% +52.5% +13.2%
Overall, from the test performance we see the following takeaways: (1) SPO can provide a strong solution very quickly, (2) in many instances, the SPO performance can be improved considerably and (3) which method yields the best improvement can vary significantly.
@@ 5.2.3 Other performance metrics and scalability
To better understand the performance of the different algorithms, in this section we provide more details on their execution.
Both the exact method (EXA) and the penalized method (PEN) often reach the time limit; in Table 6 we report the final gap values reported by Gurobi. The penalized method, in particular, shows a tendency to return exceedingly large optimality gaps. This occurs because PEN sometimes rejects the initial feasible solution and struggles to find either good bounds or high-quality feasible solutions. In one extreme case, the penalized method failed to identify a feasible solution for the shortest path instance entirely. Also note that the bipartite matching instances always finish with larger gaps than their shortest path counterpart.
@@
We also performed experiments on larger instances: with $N=1000$. Based on the analysis above, we exclude from further consideration the methods involving the exact method (EXA) and the penalized method (PEN), since their performance is dramatically affected. We report our results in Tables 7 and 8. From these results, we see that our methods remain competitive, improving the performance of SPO in many cases. In one extreme case, we obtained an improvement of 25.9% in the case of the shortest path and 39.1% in the case of bipartite matching. However, overall, we note that these improvements start to become more modest than in the smaller instances. As before, the case of bipartite matching is harder to improve than shortest path.
@@
From Table 9 we note that the number of iterations reduces dramatically as $N$ increases: this can partially explain why we observed more moderate improvements with respect to SPO in $N=1000$. We also note that, even if running local search reduces the number of iterations that ALT can perform (since the budget time is shared), the results in Tables 7 and 8 suggest that it may be worth running them in tandem.
We believe that these results suggest that a batch version of the alternating method can be a worthwhile for large instances. We strongly believe that this, along with other computational enhancements, can scale the strong result we observer in small instances.
@@ 6 Conclusions
In this work, we present an in-depth analysis of the optimization problem behind the training task of decision-focused learning. We improve on known complexity results by showing that producing regret-minimizing models is NP-complete even when only one observation is present. On the flip side, our proof of membership in NP indicates that this problem is not higher in the computational complexity hierarchy, unless the latter collapses. Additionally, we derive a non-convex quadratic optimization reformulation of the problem, whose structure we exploit empirically. Furthermore, by leveraging intermediate steps of the reformulation, we develop algorithms to address the problem at scale, including a local search procedure and an alternating direction method. These two algorithms only require solving linear programs at each iteration. When performed in tandem, they provide predictions with strong performance on both training and test sets for challenging shortest path instances and maximum weight matching problems with unknown cost/weight vectors.
@@
Acknowledgments The authors would like to thank Paul Grigas for helpful comments on this work. Víctor Bucarey and Sophia Calderón were funded by the Fondecyt project number 11220864 and the ECOS project 230033. Gonzalo Muñoz was funded by Fondecyt project number 1231522. This work was also funded by the Chilean Agency for Research and Development (ANID) through PIA/PUENTE AFB230002. Frédéric Semet was partially funded by the ECOS project C23E08.

### v5

@@ 5.2 Computational results
In this section, we compare the decision-focused predictions obtained using the aforementioned methods by evaluating the regrets they achieve. To ensure the regrets are displayed on the same scale, and following the approach in elmachtoub2022smart , we use the normalized regret instead of reporting the direct regret (as defined in (2)). The normalized regret is defined as
@@ 5.2.1 Training set performance
Figure 3 summarizes the performance of the methods on the training set for small instances of shortest paths (top) and maximum weight bipartite matching (bottom). We use the regret returned by SPO (11) as the baseline, represented as the horizontal line at zero. Each bar displays the percentage decrease/increase in regret achieved by our methods. The method with the best performance, in terms of normalized regret, is the one with the most negative value.
In the shortest path instances, the alternating method, either by itself or in tandem with local search, achieves the best performance in terms of regret. We also observe that the penalized method may often improve (decrease) the regret, but in some instances, it can dramatically increase the regret (the bars with positive value). This occurs whenever the starting point is not valid for the slice given by the chosen penalty factor in (12), causing the solver to reject the solution. For bipartite matching, there are cases where the penalized method was the best. However, the alternating method obtained the overall best results.
We also note that, in some cases, the performance of the alternating method varies considerably depending on its starting point. In most cases, starting from the solution provided by the local search algorithm yields a better outcome than starting with SPO. However, overall, the results for the alternating method are fairly robust.
We believe that these results for the training set are highly encouraging. All methods we develop here are tailored for improving the training performance in moderate running times, to which we succeeded for these challenging instances. In Appendix C, we show the detailed values for Figure 3.
In what follows, we discuss two remaining computational aspects: test set performance and scalability.
Figure 3: Training set performance on small shortest path (top) and bipartite matching (bottom) instances. Each bar represents the reduction/increment of normalized regret over the baseline SPO+.
@@ 5.2.2 Test set performance
In Figure 4 (top), we present results for shortest path instances. We observe that in 14 out of 16 cases, one of the methods we propose here improves the SPO performance. This means that at least one bar for each instance configuration is below 0. The two variants of the alternating method, which have the best in-training performance, have the best performance in 8 out of 18 test instances; however, they can increase the test regret in comparison to SPO in some cases. The results for bipartite matching, which we show in Figure 4 (bottom), are slightly different. Here, the alternating method, in its various implementations, achieves the best performance in 9 out of 18 instances –one more than in the shortest path instances. On the other hand, in 6 instances, no method was able to improve the test performance of SPO. In Appendix C, we show the detailed values for Figure 3.
Figure 4: Test set performance on small shortest path (top) and bipartite matching (bottom) instances. Each bar represents the reduction/increment of normalized regret over the baseline SPO+.
Overall, from the test performance, we see the following takeaways: (1) SPO can provide a strong solution very quickly, (2) in many instances, the SPO performance can be improved considerably, and (3) which method yields the best improvement can vary significantly.
@@ 5.2.3 Other performance metrics and scalability
To better understand the performance of the different algorithms, in this section, we provide more details on their execution.
Both the exact method (EXA) and the penalized method (PEN) often reach the time limit; in Table 1 we report the final gap values reported by Gurobi. The penalized method, in particular, shows a tendency to return exceedingly large optimality gaps. This occurs because PEN sometimes rejects the initial feasible solution and struggles to find either good bounds or high-quality feasible solutions. In one extreme case, the penalized method failed to identify a feasible solution for the shortest path instance entirely. Also, note that the bipartite matching instances always finish with larger gaps than their shortest path counterpart.
@@
We also performed experiments on larger instances: with $N=1000$. Based on the analysis above, we exclude from further consideration the methods involving the exact method (EXA) and the penalized method (PEN), since their performance is dramatically affected by instance size. We report our results in Tables 2 and 3. From these results, we see that our methods remain competitive, improving the performance of SPO in many cases. In two extreme cases, we obtained improvements of 25.9% (for shortest path) and 39.1% (for bipartite matching). However, overall, we note that these improvements start to become more modest than in the smaller instances. As before, the case of bipartite matching is harder to improve than the shortest path.
@@
From Table 4, we note that the number of iterations reduces dramatically as $N$ increases: this can partially explain why we observed more moderate improvements with respect to SPO in $N=1000$. We also note that, even if running local search reduces the number of iterations that ALT can perform (since the budget time is shared), the results in Tables 2 and 3 suggest that it may be worth running them in tandem.
We believe that these results suggest that a batch version of the alternating method can be worthwhile for large instances. We strongly believe that this, along with other computational enhancements, can scale the strong result we observed in small instances.
@@ 6 Conclusions
In this work, we present an in-depth analysis of the optimization problem behind the training task of decision-focused learning. Our proof of membership in NP indicates that this problem is not higher in the computational complexity hierarchy, unless the latter collapses. In addition, we show that the problem of determining if regret zero is achievable or not is polynomial-time solvable under mild assumptions. Additionally, we derive a non-convex quadratic optimization reformulation of the problem, whose structure we exploit empirically. Furthermore, by leveraging intermediate steps of the reformulation, we develop algorithms to address the problem, including a local search procedure and an alternating direction method. These two algorithms only require solving linear programs at each iteration. When performed in tandem, they can provide predictions with strong performance on both training and test sets for challenging shortest path instances and maximum weight matching problems with unknown cost/weight vectors.
@@
Acknowledgments
The authors would like to thank Paul Grigas for helpful comments on this work.
Appendix A: Proof of Corollary 3
Corollary 3: Consider (4) when $m(\omega,x)=\omega x$. The regret function (2) only has a finite number of values, and, furthermore, the minimum regret (4) is always attained.
Proof.
Let $I^{i}$ be an arbitrary set of indices of active constraints defining a face $F^{i}$ of $V$ for the $i$-th follower. We can consider the following system, which is similar to (6):
$(\rho^{i})^{\top}A$ $=\omega x^{i}$ $\forall i\in[N]$ (16a)
$\rho^{i}_{j}$ $>0$ $\forall\,j\in I^{i},\,\forall i\in[N]$ (16b)
$\rho^{i}_{j}$ $=0$ $\forall\,j\not\in I^{i},\,\forall i\in[N]$ (16c)
If (16) is infeasible, there is no $\omega$ “consistent” with those faces. And if (16) is feasible, every $\omega$ that is valid for (16) satisfies $V^{*}(\omega x^{i})=F^{i}$, thus the regret is the same for all of them. Since the number of possible $(I^{i})_{i\in[N]}$ is finite, we conclude. ∎
Appendix B: Proof of Theorem 2
Theorem 2: If the input for SIMPLE-REGRET is restricted to $M=0$ (i.e., determining if there is a solution with zero regret) and the data $(c^{i},x^{i})_{i=1}^{N}$ and the polytope $V$ satisfy Assumption 7, then the problem can be solved in polynomial time.
Proof.
This proof uses similar concepts to the proof of Theorem 2, but there are some important differences. We begin by noting that, since $v^{i}\in V$, it always holds that $c^{i}{}^{\top}v^{i}-z^{*}(c^{i})\geq 0$, i.e., the regret is always nonnegative. Thus, for $M=0$, SIMPLE-REGRET outputs ‘Yes’ if and only if there exists $\omega$ such that
$\max_{v^{i}\in V^{*}(\omega x^{i})}c^{i}{}^{\top}v^{i}-z^{*}(c^{i})=0\quad\forall i\in[N]$ (17)
In other words, the optimal face of
$\max_{v^{i}\in V^{*}(\omega x^{i})}c^{i}{}^{\top}v^{i}$
is contained in the optimal face defining $z^{*}(c^{i})$, i.e., $V^{*}(c^{i})$. By Assumption 7, $V^{*}(c^{i})$ is a singleton.
Let $I^{i}$ be the indices of every active constraint at $V^{*}(c^{i})$ (which can be computed in polynomial time). We claim that (17) holds for some $\tilde{\omega}$ if and only if the following linear system (over variables $\rho,\omega$) is feasible:
$(\rho^{i})^{\top}A$ $=x^{i}\omega$ $\forall i\in[N]$ (18a)
$\rho^{i}_{j}$ $=0$ $\forall j\not\in I^{i},\,\forall i\in[N]$ (18b)
$\rho^{i}_{j}$ $\geq 1$ $\forall j\in I^{i},\,\forall i\in[N]$ (18c)
We note that the $\tilde{\omega}$ certifying (17) may or may not be the same as the $\omega$ in (18). Proving this equivalence suffices, as system (18) can be solved in polynomial time.
Suppose the system (18) is feasible, and take $(\rho,\omega)$ that satisfy it. Additionally, for each $i$, take $v^{i}$ the unique optimal solution in $V^{*}(c^{i})$. We claim that each $(v^{i},\rho^{i})$ optimize $V^{*}(\omega x^{i})$ and its dual. Indeed, primal and dual feasibility hold by construction. Complementary slackness also holds by construction, as for each $j$ either $a_{j}^{\top}v^{i}-b_{j}=0$ or $\rho^{i}_{j}=0$.
Fixing the dual solutions $\rho^{i}$, and via complementary slackness again, we see that any $\tilde{v}^{i}\in V^{*}(\omega x^{i})$, must satisfy $a_{j}^{\top}\tilde{v}^{i}-b_{j}=0$ $\forall j\in I^{i}$. This means that $V^{*}(\omega x^{i})$ is a singleton. This implies (17) using $\tilde{\omega}=\omega$.
For the other direction, suppose (17) holds for some $\omega$. This directly implies that $V^{*}(\omega x^{i})=V^{*}(c^{i})=\{v^{i}\}$, since any other vector in the polytope has a strictly larger value than $c^{i}{}^{\top}v^{i}$ by assumption. By optimality conditions for $V^{*}(\omega x^{i})$, there exists $\tilde{\rho}$ such that
$(\tilde{\rho}^{i})^{\top}A$ $=x^{i}\omega$ $\forall i\in[N]$
$\tilde{\rho}^{i}$ $\geq 0$ $\forall i\in[N]$
$\tilde{\rho}^{i}_{j}(a_{j}^{\top}v^{i}-b)$ $=0$ $\forall j,\,\forall i\in[N]$
Furthermore, we can take $\tilde{\rho}^{i}$ to be strictly complementary with $v^{i}$ (as the latter is the unique optimal solution). Thus, the following holds
$(\tilde{\rho}^{i})^{\top}A$ $=x^{i}\omega$ $\forall i\in[N]$
$\tilde{\rho}^{i}_{j}$ $>0$ $\forall j\in I^{i},\,\forall i\in[N]$
$\tilde{\rho}^{i}_{j}$ $=0$ $\forall j\not\in I^{i},\,\forall i\in[N]$
We can then rescale $\tilde{\rho}$ and $\omega$ to obtain (18). ∎
Appendix C: Numerical results of Figures 3 and 4
In Tables 5 and 6 for training and 7 and 8 for test, we present the detailed results of Figures 3 and 4, respectively. We use the regret returned by SPO (model (11)) as the baseline. The subsequent columns display the percentage decrease in regret achieved by our methods. The method with the best performance in terms of normalized regret is highlighted in color.
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.035 0.0% -14.3% -17.1% -42.9% -54.3%
50 2 0.5 0.171 0.0% -22.2% -25.1% -43.3% -43.3%
50 8 0.0 0.133 0.0% -13.5% +53.4% -19.5% -19.5%
50 8 0.5 0.381 -23.6% -45.1% +9.2% -67.5% -64.8%
50 16 0.0 1.254 -66.4% -70.6% -6.7% -82.0% -75.4%
50 16 0.5 0.99 -59.5% -64.5% +253.5% -71.2% -77.8%
100 2 0.0 0.078 0.0% 0.0% 0.0% -32.1% -32.1%
100 2 0.5 0.195 0.0% 0.0% 0.0% -37.9% -40.5%
100 8 0.0 0.399 0.0% 0.0% 0.0% -20.8% -20.8%
100 8 0.5 0.572 -29.0% -29.0% -29.0% -47.6% -44.9%
100 16 0.0 3.243 -61.2% -61.2% -23.0% -69.1% -13.8%
100 16 0.5 1.945 -15.3% -15.3% +76.1% -28.0% -5.3%
200 2 0.0 0.089 0.0% 0.0% 0.0% -20.2% -20.2%
200 2 0.5 0.232 -0.9% -0.9% -0.9% -11.2% -13.4%
200 8 0.0 0.693 -16.7% -16.7% -16.7% -25.4% -4.0%
200 8 0.5 0.621 -1.4% -1.4% -1.4% -22.1% -16.1%
200 16 0.0 2.14 -60.5% -60.5% -60.5% -61.5% -2.9%
200 16 0.5 6.194 -79.4% -79.4% -79.4% -80.4% -0.2%
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.065 0.0% -12.3% -46.2% -56.9% -56.9%
50 2 0.5 0.125 0.0% -15.2% -35.2% -46.4% -46.4%
50 8 0.0 0.344 0.0% -12.8% -24.4% -31.7% -31.7%
50 8 0.5 0.249 0.0% -15.3% -11.2% -39.4% -39.4%
50 16 0.0 0.316 -3.5% -11.4% +5.1% -52.2% -50.0%
50 16 0.5 0.344 -18.3% -32.0% -18.3% -74.1% -38.4%
100 2 0.0 0.091 0.0% -4.4% -47.3% -29.7% -29.7%
100 2 0.5 0.182 0.0% 0.0% -42.3% -27.5% -27.5%
100 8 0.0 0.391 -10.7% -10.7% -10.7% -40.2% -20.2%
100 8 0.5 0.29 0.0% -5.9% 0.0% -27.6% -27.6%
100 16 0.0 0.415 0.0% -6.7% 0.0% -19.8% -19.8%
100 16 0.5 0.536 -12.9% -12.9% -12.9% -38.6% -19.8%
200 2 0.0 0.103 0.0% 0.0% 0.0% -15.5% -15.5%
200 2 0.5 0.207 0.0% 0.0% 0.0% -11.6% -11.6%
200 8 0.0 0.361 0.0% 0.0% 0.0% -8.9% -8.9%
200 8 0.5 0.387 0.0% 0.0% 0.0% -7.5% -7.5%
200 16 0.0 0.478 -17.6% -17.6% -17.6% -28.9% -17.6%
200 16 0.5 0.589 -25.3% -25.3% -25.3% -33.8% -28.5%
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.105 0.0% -2.9% -24.8% -11.4% -4.8%
50 2 0.5 0.304 0.0% -15.1% +6.9% -6.9% -7.2%
50 8 0.0 0.45 0.0% +8.2% -34.2% -52.7% -34.4%
50 8 0.5 1.056 -3.3% +8.4% +7.4% +16.5% +5.2%
50 16 0.0 3.762 -74.9% -73.1% -76.2% -72.8% -74.9%
50 16 0.5 3.504 -9.6% -7.4% +42.9% +6.2% -13.8%
100 2 0.0 0.099 0.0% 0.0% 0.0% -1.0% -1.0%
100 2 0.5 0.227 0.0% 0.0% 0.0% +7.0% +11.0%
100 8 0.0 0.776 0.0% 0.0% 0.0% +8.9% -13.9%
100 8 0.5 0.707 -3.0% -3.0% -3.0% -24.9% +35.5%
100 16 0.0 3.882 +39.6% +39.6% -12.6% -6.1% -7.5%
100 16 0.5 13.645 +13.9% +13.9% -8.2% +14.9% +5.0%
200 2 0.0 0.099 0.0% 0.0% 0.0% 0.0% -1.0%
200 2 0.5 0.238 +11.3% +11.3% +11.3% +10.9% -5.0%
200 8 0.0 0.591 -2.7% -2.7% -2.7% +1.9% +14.7%
200 8 0.5 1.023 -23.3% -23.3% -23.3% -19.8% -2.5%
200 16 0.0 1.836 +39.9% +39.9% +39.9% +27.6% +15.0%
200 16 0.5 3.384 +2.0% +2.0% +2.0% -5.8% -12.1%
N Deg Noise SPO SPO-LS SPO-LS-EXA SPO-LS-PEN SPO-LS-ALT SPO-ALT
50 2 0.0 0.118 0.0% +39.0% -2.5% -3.4% +14.4%
50 2 0.5 0.26 0.0% 0.0% -5.8% -4.2% -10.8%
50 8 0.0 0.349 0.0% +30.4% -6.9% -6.3% -10.3%
50 8 0.5 0.474 0.0% +6.1% +4.0% +6.5% +5.3%
50 16 0.0 0.781 -17.2% -20.5% -13.3% +15.1% -17.4%
50 16 0.5 0.596 -21.8% -9.4% -21.8% -9.9% +11.4%
100 2 0.0 0.118 0.0% +51.7% -16.1% +5.9% +5.9%
100 2 0.5 0.252 0.0% 0.0% -1.2% 0.0% -0.8%
100 8 0.0 0.403 +6.2% +6.2% +6.2% +5.2% 0.0%
100 8 0.5 0.431 0.0% -1.6% 0.0% -2.6% -2.1%
100 16 0.0 0.734 0.0% 0.1% 0.0% -18.5% -18.4%
100 16 0.5 0.455 +16.9% +16.9% +16.9% +52.1% -1.3%
200 2 0.0 0.126 0.0% 0.0% 0.0% -4.0% -3.2%
200 2 0.5 0.229 0.0% 0.0% 0.0% +3.1% +2.6%
200 8 0.0 0.368 0.0% 0.0% 0.0% +6.5% +6.5%
200 8 0.5 0.401 0.0% 0.0% 0.0% +1.5% +1.5%
200 16 0.0 0.739 -6.4% -6.4% -6.4% -19.4% +0.1%
200 16 0.5 0.455 +51.2% +51.2% +51.2% +52.5% +13.2%